OpenStudy (anonymous):

In a small bag of M&Ms, there are 10 candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(you select first an orange candy and then a blue candy).

OpenStudy (anonymous):

Orange first = 9/13 Blue next = 4/12 P (both) = (9/13)*(4/12) P (both) = 3/13 which rounds of to 0.23

OpenStudy (anonymous):

How did you get 9/13 for orange and 4/12 for blue?

OpenStudy (anonymous):

10+ 4 =14 14-1 = 13 you take an orange candy 10-1 = 9 orange candies so 9/13 then.... blue candies 4/13

OpenStudy (anonymous):

do u understand me?

OpenStudy (anonymous):

Yep! Thanks :)

OpenStudy (anonymous):

thanks